I do not know how to set up this problem. Can anyone help me?
Two frames are needed with the same outside perimeter: one frame in the shape of a square and one in the shape of an equilateral triangle. Each side of the triangle is 6 centimeters longer than each side of the square. Find the dimensions of each frame.

Found 2 solutions by nerdybill, scott8148:
Answer by nerdybill(7384)

Two frames are needed with the same outside perimeter: one frame in the shape of a square and one in the shape of an equilateral triangle. Each side of the triangle is 6 centimeters longer than each side of the square. Find the dimensions of each frame.
.
Let x = length of one side of the square (in cm)
then
x+6 = length of one side of the triangle
.
Perimeter of square is 4x
perimeter of triangle is 3(x+6)
.
So:
4x = 3(x+6)
4x = 3x+18
x = 18 centimeters
and
Length of one side of triangle:
x+6 = 18+6 = 24 centimeters
.
Dimensions of square: 18 by 18 by 18 by 18 (centimeters)
Dimensions of triangle: 24 by 24 by 24 (centimeters)

Answer by scott8148(6628)

the perimeters are the same
three sides to a triangle
four sides to a square

"Each side of the triangle is 6 centimeters longer than each side of the square" ___ 4s = 3(s + 6)

4s = 3s + 18